Circular No. 7416

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/2000 J1 (FERRIS) An apparently asteroidal object reported by LONEOS (observer

W. D. Ferris, measurer B. W. Koehn; discovery observation below) that was posted on the NEO Confirmation Page has been found to have a faint 10′′ tail in p.a. 150^{o} (mag R = 19.6 in a 5′′ aperture) on CCD images taken by M. Hicks with the Kitt Peak 2.13-m reflector on May 8.3 UT. Full astrometry and preliminary parabolic orbital elements appear on MPEC 2000-J30 (T = 2000 May 17.8 TT, q = 2.545 AU, i = 98^{o}.8).

2000 UT  α_{2000}  δ_{2000}  m_{2}
May 4.32273  15^{h}02^{m}30.^{s}13  +38^{o}29′04′′.6  19.0

SUPERNOVA 2000cd IN ANONYMOUS GALAXY L.-G. Strolger, University of Michigan and Cerro Tololo Interameri-

can Observatory (CTIO); R. C. Smith, CTIO; M. Hamuy, University of Arizona; M. M. Phillips, Las Campanas Observatory, and the rest of the Nearby Galaxies Supernova Search Team report the discovery of a supernova (R = 18.2) on images taken with the Kitt Peak 0.9-m telescope (+ Mosaic camera) on Mar. 24 during their March supernova search campaign (cf. IAUC 7404). SN 2000cd is located at α = 10^{h}56^{m}48^{s}.69, δ = −5^{o}34′40′′.8 (equinox 2000.0), and nothing was visible at this position on reference images taken on 1999 Feb. 23 with the same telescope. No host galaxy is visible to magnitude R = 21. SN 2000cd is identified as a type-IIn supernova at an early epoch from spectra (range 400–800 nm) obtained by Strolger and Smith using the Mayall 4-m telescope (+ R-C spectrograph) on Apr. 4. The spectra exhibit a blue continuum with broad emission features as well as narrow Hα, Hβ, Hγ, and [N II] emission; these narrow emission lines yield a redshift of z = 0.047. A finder chart is posted at

http://ctios6.ctio.noao.edu/~ngss/ngss3/aster/asterfinder.html.

SUPERNOVA 2000bi IN ANONYMOUS GALAXY Further to his reports on IAUC 7398 and 7399, M. M. M. Santangelo

provides a revised position for SN 2000bi from a deep R CCD frame (limiting mag ∼ 21) taken by A. De Blasi with the Loiano 1.52-m reflector (+ BFOSC) on Apr. 20.839 UT, when the supernova was at mag 19.3: α = 7^{h}21^{m}33^{s}.52, δ = +55^{o}45′30′′.8 (equinox 2000.0). He adds that SN 2000bi seems to be fading very slowly.

2000 May 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green